# Identifying and Mitigating API Misuse in Large Language Models

### v1

@@ 5.3 Hyperparameters
In this study, we set a maximum token length to 2048, a temperature of 0, and a top-p of 0.95. The few-shot learning approach [51] is used to prompt the models across four stages: detection, classification, reasoning, and repair suggestion. Specifically, we use two shots per stage as a demonstration to ensure the output format. For each type inside the API misuse taxonomy, we sample two failure cases to teach the models.
The prompts for each stage are designed as follows: (1) Detect — “Identify any incorrect usages in the following code snippet based on one of the following categories: Intent, Hallucination, Redundancy, Missing.” The definition of each category is provided after the instructions. To prevent LLMs from hallucinating API misuse, we append “Please answer with "No Issue" if there is no obvious API misuse.” at the end of the prompt. (2) Reason — “Think step by step and explain why this API usage is incorrect and how it deviates from the intended usage.” (3) Fix — “Propose a correction that aligns with the intended functionality of the API.”
@@
The refusal rate is introduced as an additional evaluation metric inspired by [52, 53, 54]. It measures the model’s ability to avoid hallucinating an API misuse when none exists in the code snippet. This score is defined as the percentage of instances where the model correctly refuses to detect a misuse when no misuse is present in the code. The formula for the refusal rate is:
Refusal Rate=Number of Correct RejectionsTotal Number of No Misuse $\text{Refusal Rate}=\frac{\text{Number of Correct Rejections}}{\text{Total Number of No Misuse Samples}}\times 100$
TABLE IV: 10% Sample of Misuse Types for Java and Python
@@ 5.4 Dataset
We sampled 10% of code snippets from each annotated misuse type in Section 3. As shown in Table IV, the dataset includes a representative sample of API misuse types across Java and Python. Specifically, the dataset captures the following misuse types: Intent, Hallucination, Redundancy, Position, and Missing. For each type, a proportional number of samples was extracted based on their original distribution. To further enhance the evaluation of repair approaches, including Dr.Fix and other baselines, we introduced an additional 100 snippets that exactly match the reference source code. This addition allows us to test the ability of repair tools to avoid unnecessary hallucinations or misinterpretations of correct API usage. By including these reference-aligned snippets, we can measure the precision of tools in distinguishing between actual misuse and correct implementations.
@@ 5.5 Results Analysis
TABLE V: Performance of Models across Different Methods. Dr.Fix (w/o taxonomy) means that the models are not prompted with the defined API misuse taxonomy and demonstrated examples, and perform the detection based on their own understanding. Baseline denotes the LLM-based APR method where the models are prompted to generate the fixed code snippet if they detect the API misuse inside the given snippet.
@@
Qwen-Coder-2.5 Baseline 36.3 28.4 24
@@
Qwen-Coder-2.5 Baseline 55.7 73.3 21
@@
Qwen-Coder-2.5 Baseline 15.1 43.8 14
@@
Qwen-Coder-2.5 Baseline 81.2 62.3 10
@@
The performance of the models across different methods is summarized in Table V. The results highlight the significant improvement in API misuse detection and repair capabilities achieved by incorporating the Dr.Fix methodology, with and without the taxonomy. This section provides an in-depth analysis of the findings for both Java and Python across the evaluated metrics.
@@
Java Performance: For Java, both Dr.Fix and Dr.Fix (w/o taxonomy) outperform the baseline method across all metrics. In the Method misuse category, GPT-4o with Dr.Fix achieves the highest BLEU score, CodeBERTScore, and EM, demonstrating its ability to generate syntactically and semantically appropriate repairs. The intermediate version, Dr.Fix (w/o taxonomy), also shows substantial improvements, with GPT-4o achieving a BLEU score of 76.1 and an EM of 92, which further enhance to 79.4 and 96, respectively, with the full taxonomy. Similarly, Qwen2.5-Coder and Llama-3.1 show marked improvements with both versions of Dr.Fix, with BLEU scores increasing from 36.3 to 58.3 and from 30.5 to 68.9, respectively. In the Parameter misuse category, the results follow a similar trend. GPT-4o achieves the best performance with BLEU, CodeBERTScore, and EM. Dr.Fix (w/o taxonomy) notably enhances the performance of all models, with improvements in both syntactic and semantic alignment, as reflected in the BLEU and CodeBERTScore metrics, which are further boosted by the full taxonomy.
@@
Python Performance: For Python, both Dr.Fix and Dr.Fix (w/o taxonomy) demonstrate significant improvements over the baseline method. In the Method misuse category, GPT-4o achieves the highest BLEU score and CodeBERTScore, indicating superior generalization and repair accuracy. The EM score also improves from 31 to 71 with the full taxonomy, while Dr.Fix (w/o taxonomy) achieves an EM of 66. Similarly, Qwen2.5-Coder and Llama-3.1 achieve notable gains in BLEU and CodeBERTScore metrics with both versions of Dr.Fix, showcasing the effectiveness of the methodology in improving code repair suggestions. In the Parameter misuse category, all models exhibit enhanced performance with Dr.Fix. GPT-4o achieves a BLEU score of 88.6 and a CodeBERTScore of 92.4 with the full taxonomy, surpassing other models in both accuracy and precision. While the baseline method performs reasonably well in this category, both versions of Dr.Fix provide a clear advantage, ensuring that repair suggestions align more closely with developer intentions, with the full taxonomy offering the most significant improvements.
@@
Figure 10: Refusal rate comparison of Llama-3.1, Qwen-Coder-2.5, and GPT-4o with and without Dr.Fix.
@@
Refusal Rate: The refusal rate metric evaluates the models’ ability to avoid hallucinating API misuses when no misuse exists. As shown in Figure 10, Dr.Fix significantly enhances the refusal rates across all tested models. GPT-4o achieved the highest refusal rate of 84% with Dr.Fix, demonstrating its exceptional ability to differentiate between correct and incorrect API usage. Similarly, Qwen-Coder-2.5 and Llama-3.1 showed substantial improvements, achieving refusal rates of 67% and 46%, respectively, compared to their baseline configurations, which had refusal rates of 0%, 55%, and 35%, respectively. Dr.Fix (w/o taxonomy) also showed promising results, with refusal rates of 70%, 55%, and 35% for GPT-4o, Qwen-Coder-2.5, and Llama-3.1, respectively. Notably, Dr.Fix slightly decreases the refusal rate when no API misuse taxonomy is given for the model, suggesting that it effectively reduces false positives by avoiding unnecessary refusals for correct API usage. These results highlight Dr.Fix’s effectiveness in reducing hallucinated API misuses across diverse LLM architectures, with the full taxonomy version providing the most significant improvements. {findingbox} Our experimental results demonstrate that Dr.Fix significantly enhances API misuse repair across all models, achieving up to 130% improvement in exact match rates compared to baseline approaches while maintaining high code quality. The framework reduces refusal errors by 14-20% absolute percentage points across different architectures, showcasing its effectiveness as a reliable program repair solution for diverse programming languages.
@@ 6.1 Internal Validity
One potential threat to internal validity is the reliance on subjective human annotations for the initial classification of API misuse. Although two annotators with over five years of programming experience independently reviewed the samples, biases in understanding or interpretation of the API misuse may have influenced the categorization. We mitigated this by ensuring high inter-annotator agreement, as indicated by the high Cohen’s Kappa score, and resolving disagreements through detailed discussions. Another threat arises from the inherent variability in LLM outputs due to stochastic sampling. To address this, we kept the model parameters fixed across experiments, varying only the prompts as necessary. However, given resource constraints, experiments were conducted once, and results may vary under repeated evaluations. Future studies should consider multiple iterations to assess consistency and variability in repair suggestions.
@@ 6.2 External Validity
A significant threat to external validity is the generalizability of the findings to other programming languages, APIs, and LLMs. Our dataset is limited to Python and Java, focusing on specific API misuse categories. While these languages and categories are widely used and relevant, the results may not directly extend to other contexts, such as low-level languages or highly specialized APIs. Similarly, the LLMs evaluated in this study, including GPT-4o, Qwen2.5-Coder, and Llama-3.1, represent a subset of available models. The effectiveness of Dr.Fix with other LLMs or on alternative programming datasets remains uncertain. Future research should expand the scope to include additional languages, APIs, and models to validate the framework’s broader applicability.
@@ 6.3 Construct Validity
Construct validity may be threatened by the choice of evaluation metrics and dataset design. Metrics such as BLEU and CodeBERTScore, while widely used, primarily measure syntactic and semantic similarity and may not fully capture the functional correctness of the repairs. LLMs-as-Judges [55, 56] is hence a more promising direction to explore for code quality evaluation. Moreover, the inclusion of 100 snippets without API misuse to evaluate refusal rates introduces a potential bias in the dataset composition. Although this addition aims to measure the models’ ability to avoid hallucinating misuse, its proportion relative to actual misuse cases may influence the results. Future work could explore alternative metrics or evaluation methods that better assess functional correctness and developer usability.
@@ 7.1 Large Language Models for Code
LLMs have emerged as a transformative approach for code generation tasks, broadly categorized into standard language models and instruction-tuned models. Standard language models are pre-trained on raw corpora using next-token prediction, exemplified by models such as Codex [57], CodeGen [58], CodeGeeX [59], CodeT5 [60] and StarCoder [3]. Codex, with 12B parameters, and StarCoder, featuring 15.5B parameters, demonstrated significant advancements in program synthesis through training on GitHub code and additional sources. Instruction-tuned models represent the next evolution, with examples like ChatGPT employing Reinforcement Learning with Human Feedback [61]. Open-source alternatives include WizardCoder [62], which fine-tunes StarCoder using techniques like Evol-Instruct, and InstructCodeT5+ [63], which extends CodeT5 through instruction-tuned processes. These developments have enabled zero-shot capabilities in code generation tasks, demonstrating emergent behaviors that were previously unattainable while maintaining the essential citations and academic format.
@@ 7.2 Bug Study of Large-Language-Model-Generated Code
Understanding bugs in LLM-generated code is crucial for improving code generation capabilities. While prior studies have analyzed bugs in deep learning systems broadly [64, 65, 66], recent research has focused specifically on defects in LLM-generated code [67]. [68] investigated bugs in non-standalone code generated by LLMs, though their study was limited by potential data leakage issues and the use of now-outdated models. [69] addressed these limitations by evaluating LLM-generated code with a robust benchmark designed to minimize data leakage, providing a more accurate representation of model performance in practical applications. Our work builds upon these insights by focusing specifically on API misuse patterns in LLM-generated code, introducing a taxonomy of error types and proposing the Dr.Fix repair mechanism.

### v2

@@ 5.3 Hyperparameters
In this study, we set a maximum token length to 2048, a temperature of 0, and a top-p of 0.95. The few-shot learning approach [brown2020language] is used to prompt the models across four stages: detection, classification, reasoning, and repair suggestion. Specifically, we use two shots per stage as a demonstration to ensure the output format. For each type inside the API misuse taxonomy, we sample two failure cases to teach the models.
The prompts for each stage are designed as follows: (1) Detect - “Identify any incorrect usages in the following code snippet based on one of the following categories: Intent, Hallucination, Redundancy, Missing.” The definition of each category is provided after the instructions. To prevent LLMs from hallucinating API misuse, we append “Please answer with "No Issue" if there is no obvious API misuse.” at the end of the prompt. (2) Reason - “Think step by step and explain why this API usage is incorrect and how it deviates from the intended usage.” (3) Fix - “Propose a correction that aligns with the intended functionality of the API.”
@@
The refusal rate is introduced as an additional evaluation metric inspired by [zhang2023survey, xu2024rejection, zhang2024r]. It measures the model’s ability to avoid hallucinating an API misuse when none exists in the code snippet. This score is defined as the percentage of instances where the model correctly refuses to detect a misuse when no misuse is present in the code. The formula for the refusal rate is:
[…]
5.4 Previous Approaches
We use [zhang2023evaluating] as a representative baseline of learning-aided automated program repair (APR), which outperforms previous APR methods [chen2019sequencer, zhu2021syntax]. Zhang et al.[zhang2023evaluating] systematically evaluated pre-trained models, including CodeBERT and CodeT5, on repairing API- and semantic-related bugs. Their approach fine-tunes PLMs on a large-scale dataset of human-written git commit patches and evaluates repair performance primarily on Java projects.
This work is relevant as it demonstrates how supervised PLMs can be adapted for code repair. However, it has two key limitations: (1) it focuses exclusively on Java and deep learning library APIs, which restricts its applicability to other programming languages such as Python, and (2) it depends on fine-tuning closed or specialized models, limiting generalization to diverse libraries and cross-language settings. In contrast, our work investigates prompt-based, instruction-tuned LLMs (Dr.Fix) that require no fine-tuning, operate on a general API misuse taxonomy, and extend beyond a single domain.
For comparison, we replicate their best-performing baselines, CodeBERT and CodeT5, for the Java Method and Parameter tasks, and report them alongside Dr.Fix in our evaluation.
TABLE VI: 10% Sample of Misuse Types for Java and Python
@@ 5.5 Dataset
We sampled 10% of code snippets from each annotated misuse type of StarCoder and Copilot in Section 3. As shown in Table VI, the dataset includes a representative sample of API misuse types across Java and Python. Specifically, the dataset captures the following misuse types: Intent, Hallucination, Redundancy, Position, and Missing. For each type, a proportional number of samples was extracted based on their original distribution. To further enhance the evaluation of repair approaches, including Dr.Fix and other baselines, we introduced an additional 100 snippets that exactly match the reference source code. This addition allows us to test the ability of repair tools to avoid unnecessary hallucinations or misinterpretations of correct API usage. By including these reference-aligned snippets, we can measure the precision of tools in distinguishing between actual misuse and correct implementations.
@@ 5.6 Results Analysis
TABLE VII: Performance of Models across Different Methods. Dr.Fix (w/o taxonomy) means that the models are not prompted with the defined API misuse taxonomy and demonstrated examples, and perform the detection based on their own understanding. Baseline denotes the LLM-based APR method where the models are prompted to generate the fixed code snippet if they detect the API misuse inside the given snippet.
@@
CodeBERT [zhang2023evaluating] 12.4 15.8 3
CodeT5 [zhang2023evaluating] 20.1 24.7 5
@@
Qwen2.5-Coder-Instruct Baseline 36.3 28.4 24
@@
CodeBERT [zhang2023evaluating] 13.5 17.2 2
CodeT5 [zhang2023evaluating] 22.8 26.4 6
@@
Qwen2.5-Coder-Instruct Baseline 55.7 73.3 21
@@
Qwen2.5-Coder-Instruct Baseline 15.1 43.8 14
@@
Qwen2.5-Coder-Instruct Baseline 81.2 62.3 10
@@
The performance of the models across different methods is summarized in Table VII. The results highlight the significant improvement in API misuse detection and repair capabilities achieved by incorporating the Dr.Fix methodology, with and without the taxonomy. This section provides an in-depth analysis of the findings for both Java and Python across the evaluated metrics.
@@
Java Performance: For Java, both Dr.Fix and Dr.Fix (w/o taxonomy) outperform the baseline method across all metrics. In the Method misuse category, GPT-4o with Dr.Fix achieves the highest BLEU score, CodeBERTScore, and EM, demonstrating its ability to generate syntactically and semantically appropriate repairs. The intermediate version, Dr.Fix (w/o taxonomy), also shows substantial improvements, with GPT-4o achieving a BLEU score of 76.1 and an EM of 92, which further enhance to 79.4 and 96, respectively, with the full taxonomy. Similarly, Qwen2.5-Coder-Instruct and Llama-3.1 show marked improvements with both versions of Dr.Fix, with BLEU scores increasing from 36.3 to 58.3 and from 30.5 to 68.9, respectively. In the Parameter misuse category, the results follow a similar trend. GPT-4o achieves the best performance with BLEU, CodeBERTScore, and EM. Dr.Fix (w/o taxonomy) notably enhances the performance of all models, with improvements in both syntactic and semantic alignment, as reflected in the BLEU and CodeBERTScore metrics, which are further boosted by the full taxonomy.
In addition, we include Zhang et al. [zhang2023evaluating] as representative learning-based APR baselines. Their performance on both Java Method and Parameter repair is substantially lower than instruction-tuned LLMs. For instance, CodeBERT and CodeT5 achieve BLEU scores of 12.4 and 20.1 on the Method task, and 13.5 and 22.8 on the Parameter task, respectively, which are far below GPT-4o with Dr.Fix (79.4 and 82.5). This highlights the advantage of instruction-tuned LLMs, especially when guided by structured misuse taxonomies. This gap highlights that fine-tuned PLMs trained on commit-level patches are far less effective for API misuse repair than prompt-based LLMs. Moreover, these approaches lack awareness of the distinctive misuse patterns emerging in LLM-generated code, such as hallucination and intent misuses, limiting their applicability in this setting. These results further confirm that Dr.Fix offers not only relative gains over naive prompting baselines, but also a clear advantage over prior supervised APR approaches.
@@
Python Performance: For Python, both Dr.Fix and Dr.Fix (w/o taxonomy) demonstrate significant improvements over the baseline method. In the Method misuse category, GPT-4o achieves the highest BLEU score and CodeBERTScore, indicating superior generalization and repair accuracy. The EM score also improves from 31 to 71 with the full taxonomy, while Dr.Fix (w/o taxonomy) achieves an EM of 66. Similarly, Qwen2.5-Coder-Instruct and Llama-3.1 achieve notable gains in BLEU and CodeBERTScore metrics with both versions of Dr.Fix, showcasing the effectiveness of the methodology in improving code repair suggestions. In the Parameter misuse category, all models exhibit enhanced performance with Dr.Fix. GPT-4o achieves a BLEU score of 88.6 and a CodeBERTScore of 92.4 with the full taxonomy, surpassing other models in both accuracy and precision. While the baseline method performs reasonably well in this category, both versions of Dr.Fix provide a clear advantage, ensuring that repair suggestions align more closely with developer intentions, with the full taxonomy offering the most significant improvements.
@@
Figure 10: Refusal rate comparison of Llama-3.1, Qwen2.5-Coder-Instruct, and GPT-4o with and without Dr.Fix.
@@
Refusal Rate: The refusal rate metric evaluates the models’ ability to avoid hallucinating API misuses when no misuse exists. As shown in Figure 10, Dr.Fix significantly enhances the refusal rates across all tested models. GPT-4o achieved the highest refusal rate of 84% with Dr.Fix, demonstrating its exceptional ability to differentiate between correct and incorrect API usage. Similarly, Qwen2.5-Coder-Instruct and Llama-3.1 showed substantial improvements, achieving refusal rates of 67% and 46%, respectively, compared to their baseline configurations, which had refusal rates of 0%, 55%, and 35%, respectively. Dr.Fix (w/o taxonomy) also showed promising results, with refusal rates of 70%, 55%, and 35% for GPT-4o, Qwen2.5-Coder-Instruct, and Llama-3.1, respectively. Notably, Dr.Fix slightly decreases the refusal rate when no API misuse taxonomy is given for the model, suggesting that it effectively reduces false positives by avoiding unnecessary refusals for correct API usage. These results highlight Dr.Fix’s effectiveness in reducing hallucinated API misuses across diverse LLM architectures, with the full taxonomy version providing the most significant improvements. {findingbox} Our experimental results demonstrate that Dr.Fix significantly enhances API misuse repair across all models, achieving up to 130% improvement in exact match rates compared to baseline approaches while maintaining high code quality. The framework reduces refusal errors by 14-20% absolute percentage points across different architectures, showcasing its effectiveness as a reliable program repair solution for diverse programming languages.
@@ 6.1 Internal Validity
One potential threat to internal validity is the reliance on subjective human annotations for the initial classification of API misuse. Although two annotators with over five years of programming experience independently reviewed the samples, biases in understanding or interpretation of the API misuse may have influenced the categorization. We mitigated this by ensuring high inter-annotator agreement, as indicated by the high Cohen’s Kappa score, and resolving disagreements through detailed discussions. Another threat arises from the inherent variability in LLM outputs due to stochastic sampling. To address this, we kept the model parameters fixed across experiments, varying only the prompts as necessary. However, given resource constraints, experiments were conducted once, and results may vary under repeated evaluations. Future studies should consider multiple iterations to assess consistency and variability in repair suggestions. Finally, our closed-coding process to adopt taxonomy categories may have introduced subjectivity. While grounded in prior work, different coding decisions could lead to slightly different categorization. We chose closed coding because prior studies [he2023python, wei2024demystifying] on API misuse provide taxonomies that can be readily adapted with minor modifications, making them suitable for our context. This approach allows us to capture patterns relevant to LLM-generated code without reinventing categories from scratch, though it carries a risk of overlap with existing classifications. Another potential threat concerns the breadth of our API completion tasks. We mitigated this by sampling 3,000 distinct API signatures each for Python and Java from The Stack, covering overloads and version-specific variants through static analysis. While this approach ensures diversity across frequently used APIs, the topical coverage of the underlying packages is difficult to determine, since open-source repositories often mix multiple domains and lack consistent metadata. As a result, our dataset may not fully capture the distribution of APIs across specific application areas, and we leave a more fine-grained analysis of domain coverage for future work.
@@ 6.2 External Validity
A significant threat to external validity is the generalizability of the findings to other programming languages, APIs, and LLMs. Our dataset is limited to Python and Java, focusing on specific API misuse categories. While these languages and categories are widely used and relevant, the results may not directly extend to other contexts, such as low-level languages or highly specialized APIs. Similarly, the LLMs evaluated for Dr.Fix in this study, including GPT-4o, Qwen2.5-Coder-Instruct, and Llama-3.1, represent a subset of available models. The effectiveness of Dr.Fix with other LLMs or on alternative programming datasets remains uncertain. Furthermore, our comparison with prior APR baselines (e.g., CodeBERT and CodeT5) is limited to Java and deep learning APIs, following the scope of that work. Hence, relative improvements may differ in other ecosystems. Moreover, because our dataset emphasizes frequently used APIs, the results may not fully generalize to rarely used or deprecated APIs, which could exhibit distinct misuse patterns or outdated usage behaviors. Finally, while our analysis is grounded in large-scale manual annotation, this process is extremely resource-consuming and limits the breadth of models and datasets we could feasibly include. As such, our findings may not generalize to other LLM families not studied here, and we leave broader cross-model validation to future work.
@@ 6.3 Construct Validity
Construct validity may be threatened by the choice of evaluation metrics and dataset design. Metrics such as BLEU and CodeBERTScore, while widely used, primarily measure syntactic and semantic similarity and may not fully capture the functional correctness of the repairs. LLMs-as-Judges [zhuo2024ice, he2025code] is hence a more promising direction to explore for code quality evaluation. Moreover, the inclusion of 100 snippets without API misuse to evaluate refusal rates introduces a potential bias in the dataset composition. Although this addition aims to measure the models’ ability to avoid hallucinating misuse, its proportion relative to actual misuse cases may influence the results. Finally, our use of human-written code from The Stack as the ground truth may embed biases or errors, and multiple valid API usages could be penalized under our evaluation setup. We also rely on the open-source community’s sustained effort to ensure the quality and relevance of The Stack, which may introduce variability beyond our control. In addition, the EM metric, while capturing structural matches at the AST level, may conflate syntactic similarity with semantic correctness, limiting its interpretability. Future work could explore alternative metrics or evaluation methods that better assess functional correctness and developer usability.
@@ 7.1 Large Language Models for Code
LLMs have emerged as a transformative approach for code generation tasks, broadly categorized into standard language models and instruction-tuned models. Standard language models are pre-trained on raw corpora using next-token prediction, exemplified by models such as Codex [chen2021evaluating], CodeGen [nijkampcodegen], CodeGeeX [zheng2023codegeex], CodeT5 [wang2021codet5] and StarCoder [li2023starcoder]. Codex, with 12B parameters, and StarCoder, featuring 15.5B parameters, demonstrated significant advancements in program synthesis through training on GitHub code and additional sources. Instruction-tuned models represent the next evolution, with examples like ChatGPT employing Reinforcement Learning with Human Feedback [ouyang2022training]. Open-source alternatives include WizardCoder [luowizardcoder], which fine-tunes StarCoder using techniques like Evol-Instruct, and InstructCodeT5+ [wang2023codet5+], which extends CodeT5 through instruction-tuned processes. These developments have enabled zero-shot capabilities in code generation tasks, demonstrating emergent behaviors that were previously unattainable while maintaining the essential citations and academic format.
@@ 7.2 Bug Study of Large-Language-Model-Generated Code
Understanding bugs in LLM-generated code is crucial for improving code generation capabilities. While prior studies have analyzed bugs in deep learning systems broadly [islam2019comprehensive, wang2022empirical, zhang2018empirical], recent research has focused specifically on defects in LLM-generated code [liu2024refining]. [tambon2024bugs] investigated bugs in non-standalone code generated by LLMs, though their study was limited by potential data leakage issues and the use of now-outdated models. [dou2024s] addressed these limitations by evaluating LLM-generated code with a robust benchmark designed to minimize data leakage, providing a more accurate representation of model performance in practical applications. Our work builds upon these insights by focusing specifically on API misuse patterns in LLM-generated code, introducing a taxonomy of error types and proposing the Dr.Fix repair mechanism.